A water heater has a resistance of 5.3 ohms and takes 43.5 amp when in operation. If itis in service on the average of 2 hr per day, how many kilowatt hours of energy will be expended during a 30 day month?"


1
Expert's answer
2021-05-13T18:05:18-0400

We have given,


"R = 5.3 \Omega"

"I = 43.5A"

Power consumed, "P = I^2R = (43.5)^2 \times 5.3 = 10.02kW"


Energy consumption in one day,


"E = 10.02 \times 2 = 20.04kWhr"

Hence, the energy consumption in 30 days "= 20.04 \times 30 = 601.2kWhr"